The weight of sulphuric acid needed for dissolving 3gm magnesium carbonate is:

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
Dear student
The correct option is A 3.5 gm 
MgCO3+H2SO4MgSO4+CO2+H2O84 gm          98 gm 

84 gm MgCO3 needs = 98 gm H2SO4 
 1 gm MgCO3 needs = 98/84  gm H2SO4
3 gm MgSO4 needs H2SO4=98/84×3=3.5 gm.  
thanks
team askiitians